Find a particular solution of y''+y=sinx, using the method of undetermined coefficients

1
Expert's answer
2022-02-10T15:22:15-0500


From "\displaystyle
y^{\prime\prime}+y=\sin x", the auxiliary equation is;

"m^2+1=0\\
\Rightarrow m=\pm i"

Thus, the complementary function is;

"\displaystyle
y_c=c_1\cos x+c_2 \sin x" where "c_1" and "c_2" are arbitrary constant.

Next, to obtain the particular integral using method of undetermined coefficients.

Let "\displaystyle
y_p=ax\cos x+bx\sin x", then

"\displaystyle
y_p^{\prime\prime}=2b\cos x-2a\sin x-ax\cos x-bx\cos x"

Substituting "y_p" and "y_p^{\prime\prime}" into the given DE yields;

"\displaystyle
(2b\cos x-2a\sin x-ax\cos x-bx\cos x)+(ax\cos x+bx\sin x)=\sin x\\
\Rightarrow 2b\cos x-2a\sin x=\sin x"

comparing yields;

"\displaystyle
b=0," and "\displaystyle
a=-\frac{1}{2}"

Thus, "\displaystyle
y_p=ax\cos x+bx\sin x=-\frac{x\cos x}{2}+0\times x\sin x=-\frac{x\cos x}{2}"


Hence, the general solution is;

"\displaystyle
y=y_c+y_p=c_1\cos x+c_2 \sin x-\frac{x\cos x}{2}"